### v2

# ROOFS: RObust biOmarker Feature Selection

Abstract
Feature selection (FS) is essential for biomarker discovery and clinical predictive modeling. Over the past decades, methodological literature on FS has become rich and mature, offering a wide spectrum of algorithmic approaches. However, much of this methodological progress has not fully translated into applied biomedical research. Moreover, challenges inherent in biomedical data, such as high-dimensional feature space, low sample size, multicollinearity, and missing values, make FS non-trivial. To help bridge this gap between methodological development and practical application, we propose ROOFS (RObust biOmarker Feature Selection), a Python package available at https://gitlab.inria.fr/compo/roofs, designed to help researchers in the choice of FS method adapted to their problem. ROOFS benchmarks multiple FS methods on the user’s data and generates reports summarizing a comprehensive set of evaluation metrics, including downstream predictive performance estimated using optimism correction, stability, robustness of individual features, and true positive and false positive rates assessed on semi-synthetic data with a simulated outcome. We demonstrate the utility of ROOFS on data from the PIONeeR clinical trial, aimed at identifying predictors of resistance to anti-PD-(L)1 immunotherapy in lung cancer. Of the 34 FS methods gathered in ROOFS, we evaluated 23 in combination with 11 classifiers (253 models) and identified a filter based on the union of Benjamini-Hochberg false discovery rate-adjusted p-values from t-test and logistic regression as the optimal approach, outperforming other methods including widely used LASSO. We conclude that comprehensive benchmarking with ROOFS has the potential to improve the reproducibility of FS discoveries and increase the translational value of clinical models.
Keywords: feature selection, stability, biomarker discovery, benchmark, software
1 Introduction
Feature selection (FS) refers to a class of methods aimed at reducing the feature space by selecting a subset of size $k<p$, where $p$ denotes the number of features. In supervised learning, FS is used either for discovery, to identify the most informative variables, or for prediction, to reduce variance. Over the past decades, a wide range of FS methods has been proposed (see [20] and [34] for extensive reviews). FS methods can be classified into 4 groups: filter, embedded, wrapper, and ensemble methods. Filters evaluate features according to a specific criteria and output either a complete ranking or a subset selected using a predefined threshold (e.g., a statistical test). Embedded methods perform FS simultaneously with model learning, as in regularization-based approaches or tree-based models with constrained depth. Wrapper methods iteratively search for feature subsets that maximize the predictive performance of a given model. Ensemble methods aggregate outputs from different selection algorithms or from repeated runs of a single method on resampled data. Filter and embedded methods are typically more straightforward, computationally efficient, and interpretable compared to wrapper and ensemble techniques. However, computationally expensive and algorithmically complex wrapper methods possess an important property: an ability to capture feature interactions tailored to a specific classifier, typically resulting in modest performance improvement. The computational cost of the ensemble methods is balanced by their ability to select features with the strongest signal and avoid false discoveries.
A basic requirement for FS is that the reduced model should perform equivalently to or better than the full model, capturing the complexity of the original data. In the context of biomarker discovery and clinical predictive models, the selected subset, often called a “signature“, should meet additional requirements: (i) contain biologically or clinically interpretable features; (ii) be as minimal as possible; and (iii) demonstrate stability, i.e. the selected features should not be sensitive to small changes in the data [31]. FS stability is of key importance for ensuring reproducibility. It is typically assessed by measuring the similarity between feature subsets obtained across resamples of the original data, e.g. via bootstrapping or subsampling. A method that is not stable may result in models that do not generalize well to external populations, unlikely to be clinically useful.
Feature selection is a long-standing field of research in machine learning [20]. However, much of the methodological progress has not fully translated into applied biomedical research, where practitioners often rely on at most a limited set of familiar methods. Likewise, bootstrap-based procedures for assessing stability are rarely used, despite its importance having been frequently acknowledged [31, 23, 25]. In addition, the systematic evaluation of the suitability of FS methods for a given dataset is often omitted, although previous benchmark studies [5, 15] have shown that, in terms of predictive performances, the optimal method is contingent upon the specific dataset. Furthermore, FS performance is sensitive to hyperparameter choices, such as the selected subset size, and different settings can lead to different rankings of methods on the same dataset [48].
To address this gap between methodological and applied research, we propose a Python package called ROOFS (RObust biOmarker Feature Selection), available at https://gitlab.inria.fr/compo/roofs. It conducts comprehensive benchmarking of multiple pre-implemented FS algorithms on the user’s data based on three criteria: 1) method stability and robustness of individual features, 2) predictive performances, and 3) true positive and false positive rates (from semi-synthetic data with a simulated outcome). For point 2), we relied on the optimism-correction (OC) framework emphasized as a guideline for the development of clinical predictive models by Collins et al. [14]. ROOFS provides automated reporting of the evaluation metrics listed above – a key feature of the package aimed to facilitate comparison and choice of an FS method adapted to the user’s problem.
We demonstrate the practical value of ROOFS by applying it to a complex dataset from the PIONeeR clinical study (NCT03833440, Precision Immuno-Oncology for Advanced Non-small Cell Lung Cancer Patients With PD-1 ICI Resistance) to predict primary resistance to immunotherapy. This dataset is an interesting and challenging use case for FS because: 1) the data is rich and heterogeneous combining clinical features, routine blood tests and detailed immunoprofiling of blood and tumor samples measured by different clinical and research partners; 2) it contains clusters of highly correlated features, e.g. numerous variables describing related immune cell populations; and 3) $p$ (374) is roughly equivalent to $n$ (435). This is different from the three types of datasets popular in the literature: a) high-dimensional omics ($p\gg n$); b) data with a sufficiently high event-per-variable ratio, e.g. the number of patients in the smallest class is at least ten times the number of predictors; and c) big data (e.g., from electronic health records), with $n\gg p$ typically modeled using deep artificial neural networks.
A comprehensive survey of FS benchmarks published before 2020 across multiple domains is provided by Hopf and Reifenrath [26]. Table 1 summarizes more recent benchmarks on high-dimensional biomedical data. Most of these studies focused on a single data modality (genomic data). Only two studies included FS methods representing all algorithmic families, and stability was evaluated in only three of the eight benchmarks. The novelty of our study are that: 1) it addresses heterogeneous, mixed-type data; 2) it evaluates a broader and more diverse range of FS algorithms; 3) it reports all standard performance metrics, with an emphasis on stability; and 4) it extends FS evaluation to assess the balance between true and false discoveries.
Reference Data type #D #M FS Methods Metrics
F E W Ens Predictive performance Stability Runtime
Cattelani et al. [10] Genomics 8 7
Demircioğlu [15] Radiomics 50 9
Labory et al. [33] Omics 3 5
Budhraja et al. [7] Omics 4 11
Li et al. [35] Cancer multi-omics 15 8
Bommert et al. [6] Gene expression survival data 11 14
Bhadra et al. [4] Cancer multi-omics 5 5
Chen et al. [11] Not specified 4 12
Our work Cancer multi-type 1 23
#D = number of datasets, #M = number of FS methods, F = filters, E = embedded methods, W = wrappers, Ens = ensemble methods
Table 1: Recent feature selection benchmarks on high-dimensional biomedical data
2 Methods
2.1 Data
2.1.1 PIONeeR dataset
The development of ROOFS was motivated by the following real-world problem: identifying robust biomarkers of resistance to immunotherapy in lung cancer. To this end, we relied on the PIONeeR clinical trial dataset, derived from a prospective, multicenter cohort of patients (17 centers in France) with advanced or recurrent non-small cell lung cancer (NSCLC) [2]. Patients received either frontline combination therapy with platinum-based chemotherapy plus anti-PD-(L)1 immune checkpoint inhibitors (ICIs) or second-/third-line monotherapy with anti-PD-(L)1 ICIs following progression on prior platinum-based chemotherapy. The dataset integrated multi-modal data from $n=435$ patients, comprising a total of $p=374$ variables. It included clinical and demographic variables (e.g., sex, age, $p=10$), routine blood tests (e.g., blood counts, biochemistry, $p=49$), tumor multiplex immunohistochemistry markers (e.g., immune cells, $p=159$), and circulating immune and vasculophenotyping markers. The latter included both flow cytometry data ($p=141$) and soluble ($p=15$) markers. The dataset contained 30.8% of missing values. The outcome to predict was primary resistance to immunotherapy, defined as disease progression within 6 months of treatment [45].
2.1.2 Reduced PIONeeR dataset
To address multicollinearity, a reduced dataset with 214 features was derived from the full data by applying iterative variance inflation factor (VIF) pre-filtering [28], sequentially removing the feature with the highest VIF until all remaining features had VIF $\leq$ 5. At each iteration, VIF of the $j_{\text{th}}$ feature was computed as:
$\text{VIF}_{j}=\frac{1}{1-R_{j}^{2}}$
where $R_{j}^{2}$ stands for the coefficient of determination from regressing the $j_{\text{th}}$ feature onto all remaining predictors.
2.1.3 Simulated outcome
To evaluate each method’s ability to identify the "true features" (i.e., the ones hypothesized as having generated the outcome), a part of FS benchmarking implemented in ROOFS is conducted on a semi-synthetic dataset with a simulated outcome denoted $\tilde{y}$. With the true model being unknown, the real outcome $y$ cannot be used directly for this purpose, but the original $n\times p$ features matrix X can still be used. The synthetic outcome $\tilde{y}^{(i)}$ for observation $i$ is generated by applying a logistic function to a linear combination of $S$ pre-selected features from the real dataset X with added Gaussian noise:
$\text{score}^{i}=\frac{1}{1+\exp\left(-\Big(\sum_{j\in\mathcal{S}}\beta_{j}X_{j}^{i}+\varepsilon^{i}\Big)\right)},\quad\varepsilon^{i}\sim\mathcal{N}(0,1)$
[…]
where $\mathcal{S}$ is the index set of the pre-selected true predictors and $\beta_{j}$ is the coefficient of the $j^{th}$ predictor. The coefficients are determined using one of the two approaches: 1) Random sampling (default approach in ROOFS): sample from $N(\mu,\sigma^{2})$ and randomly assign the sign (the default parameters $\mu=2$ and $\sigma=0.5$); 2) Estimation from the data: fit an unpenalized logistic regression on the true outcome $y$ with this set of predictors and set $\beta_{j}$ equal to the estimated coefficient for predictor $j$.
For the benchmark on PIONeeR data, 19 true predictors were selected from the full dataset using a 0.01 threshold on Benjamini-Hochberg adjusted p-values from a multivariable Cox model for progression-free survival. This approach was chosen to mimic a realistic data structure. To preserve the relative effect magnitudes of the true predictors, we used approach (2) above to derive the coefficients.
The approach outlined above uses a linear model for simplicity, but ROOFS also supports nonlinearity through polynomial power terms and pairwise feature interactions.
Figure 1: Overview of the ROOFS pipeline for comprehensive FS benchmarking.
2.2 Suggested framework
The evaluation of a given FS method by ROOFS is based on a bootstrap experiment (Figure 1). Let $\mathcal{D}_{\text{original}}=\{(x^{i},y^{i})\}_{i=1}^{n}$ denote the original dataset provided by the user. From this, $B$ bootstrap resamples are derived, each denoted $\mathcal{D}_{b}$, $b\in[1,B]$. In the benchmark on PIONeeR, the number of resamples was set to 100. An identical pipeline, including data preprocessing (imputation using the median for continuous variables or mode for categorical variables and z-score normalization), FS and subsequent predictive modeling, is applied to both the full data and each bootstrap, and repeated across all user-specified combinations of FS methods and classifiers.
2.2.1 FS methods
ROOFS gathers 34 FS methods from all algorithmic families – filters, embedded methods, wrappers, and ensembles – including both classical statistical techniques and recent ML-based approaches. Most implementations were adapted from existing Python libraries, including scikit-learn [42], scikit-feature [34], Stabl [24], and others.
Of the 34 FS methods implemented in ROOFS, 23 representative approaches were included in the present benchmark. Among embedded methods, we evaluated LASSO [49] and its variants (adaptive LASSO [55] and exclusive LASSO [54]), as well as ensemble bootstrap-based methods designed to enhance stability and control false discoveries (modified Bolasso [1] with a frequency threshold of 0.5 followed by a second LASSO fit, stability selection [37], Stabl LASSO [24], RENT [29], and HSIC LASSO [51]). Both univariate and multivariate filters were used, including information theory-based techniques (CIFE [36], CMIM [19], DISR [38], JMI [52]) and similarity-based approaches (ReliefF [46], Fisher score [34]), correlation-based clustering (hierarchical clustering), and statistical methods (Gini index, t-score, minimum redundancy maximum relevance (mRMR) [34], and adjusted p-values using Benjamini-Hochberg method [3]). Wrappers included two recursive elimination methods (LR-RFE, based on logistic regression coefficients, and RF-RFE, based on random forest importances), a forward selection method (forward RF), and the SHAPley-based Shapicant [8]. Finally, random feature selection was included as a control. A complete description of all benchmarked methods is provided in Supplementary Table LABEL:sup_tab1.
2.2.2 Classifiers
ROOFS evaluates the predictive performance of signatures resulting from the FS step by training downstream classifiers on the corresponding post-selection datasets. ROOFS allows using any scikit-learn-compatible classifier. Since predictive performances can depend on the interactions between the FS method and the classifier, it is recommended to test a wide range of modelling approaches, such as linear models, tree-based methods, and ensembles. The classifiers used in the present benchmark study are listed in Supplementary Table S2.
2.2.3 Optimism correction
ROOFS uses optimism correction (OC) for evaluation of the predictive performances, which allows not to lose part of the data due to train/test splitting [14]. Three OC methods were implemented: Harrell [22], .632 [18], and .632+ [17]. The following set of metrics are calculated: AUC, accuracy, sensitivity, specificity, positive predictive value, negative predictive value.
In Harrell’s classical method, optimism is defined as the average difference between model performance on bootstrap samples (in training) and on the original dataset (in testing):
$O=\frac{1}{B}\sum_{b=1}^{B}\left(\theta_{\text{boot}}^{(b)}-\theta_{\text{orig}}^{(b)}\right),$
where $\theta_{\text{boot}}^{(b)}$ is the training performance of the $b$-th bootstrap model, $\theta_{\text{orig}}^{(b)}$ is the test performance of the same model on the original dataset, and $B$ is the number of bootstrap replicates.
The final optimism-corrected performance is obtained by subtracting the average optimism from the apparent performance (i.e., the training performance on the full dataset):
$\theta_{\text{Harrell}}=\theta_{\text{app}}-O,$
where $\theta_{\text{app}}$ denotes the apparent performance.
The .632 method accounts for the overlap between the bootstrap samples and the original data and computes $\theta_{\text{orig}}$ only on out-of-bag (OOB) samples (not included into the bootstrap, therefore not seen during training). The performance estimate is computed as
$\theta_{.632}=0.368\cdot\theta_{\text{app}}+0.632\cdot\theta_{\text{out}},$
where $\theta_{\text{out}}=\frac{1}{B}\sum_{b=1}^{B}\theta_{\text{OOB}}$. The weight $0.632$ corresponds to the approximate proportion of samples included in a bootstrap [12].
The .632+ method improves the .632 rule by modifying the weights in order to take into account the amount of overfitting:
$\theta_{\text{.632+}}=(1-w)\cdot\theta_{\text{app}}+w\cdot\theta_{\text{out}},$
where
$w=\frac{0.632}{1-0.368\cdot R}$
and $R$ is the overfitting ratio computed as:
$R=\frac{\theta_{\text{app}}-\theta_{\text{out}}}{\theta_{\text{app}}-\gamma},$
where $\gamma$ corresponds to the predictive performance on the original dataset when the outcome is randomly permuted. When there is no overfitting (i.e., the apparent score is the same as the out-of-bag score), $R$ equals to 0, $w$ equals to 0.632, and the .632+ rule replicates the .632 rule. In the presence of overfitting (apparent score is greater than the out-of-bag score), $R\in(0,1]$ and $w\in(0.632,1]$. In a setting of complete overfitting, $w$ equals to 1 and $\theta_{\text{.632+}}=\theta_{\text{out}}$. The results reported in this paper were obtained using .632+ method as it has been demonstrated to have the lowest bias in multiple settings [27].
2.2.4 FS stability
The FS stability in ROOFS is calculated from the bootstrap-derived feature sets using Nogueira’s frequency-based measure (see [41] for details and a review of alternative stability metrics):
$S=1-\frac{\frac{1}{p}\sum_{f=1}^{p}s_{f}^{2}}{\frac{\overline{k}}{p}(1-\frac{\overline{k}}{p})}$
where $s_{f}^{2}=\frac{B}{B-1}\hat{p}_{f}(1-\hat{p}_{f})$ is the unbiased sample variance of the frequency of selection $\hat{p}_{f}$ of $f^{th}$ feature over $B$ bootstraps and $\overline{k}$ is an average size of the selected subsets. $S$ takes values from $\frac{1}{B-1}$ to 1, with the lower bound asymptotically approaching 0 as $B$ approaches infinity. $S=1$ indicates perfect stability when the same features are selected at each iteration. $S=0$ corresponds to the case of random selection.
Additionally, ROOFS reports simple measures of individual feature robustness. For each feature, it computes the selection frequency obtained by every FS method and identifies how many methods select that feature with high frequency (as defined by the user; we used a 50% threshold). The selection frequency serves as an indicator of the signal strength: features that are weakly associated with the outcome appear in only a small fraction of bootstrap models, whereas features with a strong signal are consistently selected in a large proportion of bootstrap models and across multiple FS methods. This enables to distinguish between features that are stable and predictive and those that may be artifacts of a specific algorithm or over- or undersampling of particular patient subclusters.
2.2.5 Trade-off between true and false discoveries
ROOFS additionally benchmarks FS methods on a semi-synthetic dataset with simulated outcome $\mathcal{D}_{\text{semi-synthetic}}=\{(x^{i},\tilde{y}^{i})\}_{i=1}^{n},\quad x^{i}\in X,\,\tilde{y}^{i}\in\tilde{Y}$. The comparison of FS methods on a semi-synthetic data allows assessing the trade-off between true and false discoveries, since the underlying true model is known. The true positive rate (TPR) – the proportion of true predictors correctly identified – represents the discovery power of an FS algorithm. The false positive rate (FPR) – the proportion of features not included in the true model but selected by an algorithm – shows the cost of discovery in terms of error. The procedure used to derive the synthetic outcome in our benchmark on PIONeeR dataset is described in subsection 2.1.3. The variability in FPR and TPR is assessed by repeating the experiment with resampled noise, with the number of replicates specified by the user (set to 40 in our benchmark).
FS methods gathered in ROOFS can be subdivided into two groups based on how they determine the number of features they select: 1) methods with fixed, user-defined subset size, and 2) methods with varying subset size dependent on internal criteria, such as cross-validated lambda in the case of LASSO or p-value thresholds in the case of p.adjust or Shapicant. In the experiment on the semi-synthetic dataset, methods from the first group were configured to select 19 features, corresponding to the number of predictors in the true model. This configuration provided these methods with an advantage compared to the second group of methods with respect to both true and false positives. Their TPR was lower bounded, as by design the selection conducted by these methods was not too stringent, and only a random or very inaccurate method would have obtained a low TPR. In turn, their FPR was upper bounded, as these methods could not select more than 19 features.
2.2.6 Automated reporting
To allow users to efficiently evaluate the results of benchmarking, ROOFS generates a report summarizing the key metrics of FS performance. This report includes: 1) a table that aggregates FS frequencies across methods and shows agreement between different methods in selecting individual features; 2) a stability – AUC plot for a quick visual overview of performance and identification of Pareto-optimal methods; 3) a summary table that combines all metrics from every model, with filtering options that help identify the optimal methods based on the user’s criteria of interest. For an example, see the PIONeeR FS report generated with ROOFS: https://compo.inria.fr/pioneer-website/technical_reports/ML_supervised/feature_selection/main_study/
2.2.7 Choice of FS method
ROOFS allows users to manually choose a method that best fits their objectives based on the metrics included in the automatically generated report. By itself, the package does not select the best FS method from user’s data, as the acceptable trade-off between different metrics depends on the specific prediction or discovery goals.
2.2.8 Experimental setup
All experiments on PIONeeR data were conducted on a MacBook Pro (2019) with a 2.4 GHz 8-core Intel Core i9 processor and 64 GB 2667 MHz DDR4 RAM. The analysis was performed with Python 3.10.2 and R 4.2.2.
3 Results
Figure 2: Instability of LASSO on the full PIONeeR dataset ($p=374$). A: Bootstrap selection frequencies for features selected by LASSO in at least 1 bootstrap sample. B: Variability in selected subset sizes and out-of-bag AUC.
3.1 LASSO-derived signature shows modest stability
Motivated by the goal of establishing a biomarker signature capable of explaining resistance to immunotherapy in advanced lung cancer patients from a PIONeeR study, we first applied our benchmarking pipeline, as described in Methods and illustrated in Figure 1, using classical LASSO (Least Absolute Shrinkage and Selection Operator) as a baseline FS method. LASSO remains a standard FS approach in biomedical data analysis due to its simplicity of interpretation and computational efficiency, but it has been shown to be inconsistent (i.e., unable to recover a true model) in the presence of linear dependencies among features, which is common in high-dimensional problems [53]. On the PIONeeR data, LASSO showed moderate stability (S = 0.34, Figure 2A) and considerable variability in features selected across bootstrap datasets, with 255 out of 374 features selected in at least 1 bootstrap (Figure 2A). LASSO-selected feature sets varied substantially in size, ranging from 13 to 42 predictors, and the corresponding post-selection models, trained using classical ML classifiers, achieved out-of-bag AUC values ($\theta_{\text{out}}$ in the Methods) between 0.49 and 0.78. The apparent LASSO model selected 25 features, of which only 12 were selected with high confidence (defined here as being chosen in at least 50% of bootstrap models), and none were selected across all bootstraps. Low selection frequency of the remaining 13 features raised concerns about their potential clinical relevance and generalizability of the signature.
3.2 Iterative VIF increases FS stability
Since LASSO instability could be partially attributed to multicollinearity among predictors, we next applied iterative VIF pre-filtering to the full data (374 features) to derive a feature set with lower redundancy and evaluate its impact on FS stability. This produced a reduced dataset of 214 features (fig. 3A). The mean $R^{2}$ obtained from a regression of each feature onto all other predictors decreased from 0.91 (range: 0.41-0.99) in the full dataset to 0.63 (range: 0.22-0.79) following VIF pre-filtering.
To evaluate whether this reduction improved FS stability, we conducted two separate benchmarks of 23 FS methods on the full and reduced datasets. VIF pre-filtering led to a small increase in stability of the majority of FS methods (fig. 3B), likely driven by the fact that fewer features were used interchangeably across bootstrap models. Based on this improvement, all subsequent results are reported on the post-VIF dataset.
Figure 3: Improvement in FS stability following multicollinearity reduction with VIF pre-filtering compared to the original dataset. A: Correlation structure of the dataset before and after applying VIF pre-filtering. B: Absolute change in FS stability across representative FS methods.
Table 2: Results of FS benchmarking on the post-VIF PIONeeR dataset ($n=435,p=214$)
Method AUC Stability FPR TPR Run time (s)
LASSO 0.69 GB 0.34 0.02 ± 0.01 0.59 ± 0.09 0.27 ± 0.02
Adaptive LASSO 0.67 St 0.32 0.05 ± 0.01 0.59 ± 0.06 0.24 ± 0.04
Exclusive LASSO 0.66 St 0.37 0.06 ± 0.00 0.27 ± 0.03 0.59 ± 0.06
Bolasso 0.69 SVC 0.34 0.01 ± 0.01 0.49 ± 0.07 17.48 ± 0.71
Stability selection 0.67 St 0.3 0.01 ± 0.01 0.46 ± 0.07 8.10 ± 0.6
Stabl LASSO 0.64 St 0.25 0.00 ± 0.01 0.31 ± 0.07 47.36 ± 5.2
RENT 0.69 Bag 0.29 0.01 ± 0.01 0.47 ± 0.07 46.48 ± 4.23
HSIC LASSO 0.69 Ex 0.25 0.06 ± 0.01 0.41 ± 0.08 0.69 ± 0.06
CIFE 0.67 MLP 0.73 0.08 ± 0.00 0.16 ± 0.03 3.4 ± 0.14
CMIM 0.62 Bag 0.18 0.08 ± 0.01 0.14 ± 0.07 3.62 ± 0.15
DISR 0.61 Ex 0.54 0.08 ± 0.00 0.21 ± 0.02 10.11 ± 0.23
JMI 0.6 Bag 0.18 0.08 ± 0.00 0.21 ± 0.00 3.79 ± 0.07
ReliefF 0.68 MLP 0.49 0.06 ± 0.01 0.36 ± 0.13 0.26 ± 0.03
Fisher score 0.7 GB 0.37 0.03 ± 0.01 0.69 ± 0.06 0.19 ± 0.03
Gini index 0.69 Ex 0.32 0.04 ± 0.01 0.60 ± 0.07 1.00 ± 0.04
t-score 0.7 GB 0.36 0.03 ± 0.01 0.71 ± 0.06 0.11 ± 0.03
mRMR 0.7 LDA 0.32 0.04 ± 0.01 0.63 ± 0.07 0.82 ± 0.03
Hierarchical clustering 0.7 GB 0.23 0.06 ± 0.00 0.38 ± 0.04 9.44 ± 0.47
p.adjust t-test + logistic .05 0.72 GB 0.39 0.10 ± 0.03 0.92 ± 0.06 1.95 ± 0.09
LR-RFE 0.69 LDA 0.38 0.03 ± 0.01 0.7 ± 0.06 0.11 ± 0.03
RF-RFE 0.7 GB 0.38 0.02 ± 0.02 0.45 ± 0.11 52.27 ± 0.30
Forward RF 0.69 LR 0.34 0.02 ± 0.02 0.47 ± 0.13 16.29 ± 0.13
Shapicant 0.7 Bag 0.31 0.06 ± 0.01 0.7 ± 0.08 25.96 ± 0.20
Random FS (fixed size) 0.58 Ex 0 0.09 ± 0.00 0.05 ± 0.00
Random FS 0 0.06 ± 0.03 0.08 ± 0.04
Full data 0.71 GB 1
AUC is reported for the optimal FS method-classifier pair across all tested combinations. Values are presented as point estimates or mean ± SD (standard deviation); run times correspond to the average duration of the FS step (excluding predictive modeling) across bootstraps. Corresponds to better than LASSO. GB: Gradient boosting classifier; St: stacking classifier; Ex: extra trees classifier; Bag: bagging classifier; LR: logistic regression; SVC: support vector classifier; LDA: linear discriminant analysis; MLP: multi-layer perceptron; LASSO: least absolute shrinkage and selection operator; HSIC: Hilbert-Schmidt independence criterion; RENT: repeated elastic net technique; CIFE: conditional infomax feature extraction; CMIM: conditional mutual information maximization; DISR: double input symmetrical relevance; JMI: joint mutual information; mRMR: minimum Redundancy Maximum Relevance; LR-RFE: logistic regression-based recursive feature elimination.
3.3 Comprehensive benchmarking shows variability in FS stability and performance
On the reduced post-VIF dataset, filters were generally the most stable and predictive family of FS methods, although variability was observed within the group. Specifically, statistical filters achieved high optimism-corrected AUC values (ranging from 0.7 to 0.72), and several of them demonstrated good stability, for example, Fisher score (S = 0.37), t-score (S = 0.36), and p.adjust (S = 0.39) (see Table 2). In contrast, CIFE and DISR, filters based on information theory scores, achieved top stability values (S = 0.73 and 0.54, respectively) but lower predictive performance (optimism-corrected AUC = 0.67 and 0.61, respectively). All information theory filters (CIFE, CMIM, DISR, JMI) often omitted known biomarkers (e.g. lactate dehydrogenase and C-reactive protein), showing that selection can be consistent but wrong. CMIM and JMI were both unstable and not predictive, although still outperforming random FS. Embedded methods exhibited the lowest stability among other groups at comparable predictive performance. Surprisingly, embedded resampling-based ensembles designed to improve the consistency of FS (Bolasso, stability selection, Stabl LASSO, RENT, and HSIC LASSO) did not outperform standard LASSO neither in stability nor in AUC. Their stringent selection (with the exception of HSIC LASSO) may account for their lower predictive performance, but it does not explain the lack of improvement in stability (S = 0.34, 0.3, 0.25, 0.29, and 0.25, respectively, vs LASSO S = 0.34). The instability of these methods could be attributed to the variability behind the resampling procedures on which they rely, although this requires further investigation. The performance of the wrapper methods was intermediate, with AUC (0.69-0.70) and stability (0.32-0.38) falling between filter methods (the best) and embedded methods (the worst).
Among all methods, the p.adjust, a statistical filter that combined t-test and logistic regression adjusted p-values at a 0.05 threshold, achieved the best balance between stability (S = 0.39) and performance (AUC = 0.72) (Figure 4A). The gradient boosting (GB) model trained on the 22-feature signature selected using the p.adjust achieved predictive performance equivalent to that of the GB model trained on the full feature set (214 features) without FS (AUC = 0.71). Note that the strong performance of the FS-free GB model can be attributed to the restriction of the depth of tree-based classifiers in the benchmark to 2, which added an additional internal layer of FS.
Overall, most benchmarked methods achieved optimism-corrected AUC values comparable to standard LASSO, with statistical filters and wrappers showing modest improvements and information-theory filters performing worse. However, more than half of the methods demonstrated higher stability than LASSO, supporting the idea that it is possible to select features more consistently while maintaining similar predictive performance [41]. Across all methods, the association between stability and AUC was weak (Spearman’s $r=0.32$, $p=0.044$). The least stable approaches tended to perform poorly, but high stability did not necessarily translate into top predictive performances.
Figure 4: Performance of representative FS methods from different algorithmic families (top: FS methods with fixed, user-defined subset size; bottom: FS methods with varying subset size dependent on algorithm thresholds). Blue: embedded methods; violet-blue: resampling-based ensemble methods; green: filters; rose: wrappers. A, C: Trade-off between discovery and error in the experiment on the semi-synthetic PIONeeR dataset with a simulated outcome. The true positive rate (TPR) denotes the proportion of true predictors correctly identified by each FS method, while the false positive rate (FPR) shows the proportion of non-informative features incorrectly selected. Error bars represent one-half of the standard deviation (SD). B, D: Stability and predictive performance (AUC) in the experiment on the real post-VIF PIONeeR dataset.
3.4 Benchmarking on semi-synthetic data shows two distinct patterns in the relationship between true and false discoveries
The second part of the benchmark was conducted on a semi-synthetic dataset to assess the ability of FS methods to discover true predictors. As detailed in Methods, a synthetic outcome variable was simulated with a simple linear model from 19 pre-selected features from the real PIONeeR dataset, and FS methods were applied to the original dataset combined with this simulated outcome. Their performance was measured using true positive rate (TPR), false positive rate (FPR), false discovery rate (FDR) and false omission rate (FOR).
Even though the outcome was generated with a simple linear model, no method achieved both high TPR and low FPR (Table 2, fig. 4A, C). By construction, fixed-size methods exhibited an inverse relationship between TPR and FPR, as true discoveries constrained the number of false positives (fig. 4A). In contrast, threshold-based methods displayed the opposite behavior: methods with higher TPR also had higher FPR (fig. 4C). More relaxed methods (e.g., LASSO) showed higher TP and FP rates, whereas more stringent methods (e.g., Stable LASSO) produced lower TP and FP rates. The same behavior was observed when FPs were represented as the false discovery rate (FDR), i.e., the proportion of FPs relative to the signature size (supplementary fig. S1).
As expected, embedded ensemble methods that control for false discoveries (Bolasso, stability selection, Stabl LASSO, RENT) achieved the lowest FPR but at the cost of lower TPR. These methods are characterized by consistent strict selection of features with the strongest signal. However, they omit a high proportion of true predictors with weaker signals, demonstrating the TPR of only 0.49 ± 0.07, 0.46 ± 0.07, 0.31 ± 0.07 and 0.47 ± 0.07, respectively. Statistical filters exhibited the highest TPR values, while their FPR remained moderate, with one exception: p.adjust filter achieved the highest TPR (0.93 ± 0.06), rarely omitting true predictors, but at the cost of high FPR and FDR, despite Benjamini-Hochberg adjustment being explicitly designed to control false discoveries. FPR of p.adjust was significantly higher than that of random FS (0.1 ± 0.03 vs 0.07 ± 0.03; Mann–Whitney U test, $p<0.001$), while its FDR was high but significantly lower than that of random FS (0.53 ± 0.06 vs 0.89 ± 0.04; Mann–Whitney U test, $p<0.001$). The high FPR of p.adjust may be explained by residual correlations among predictors in the post-VIF dataset, which can inflate false discoveries in multiple hypothesis testing settings, as demonstrated by Kanduri et al. [32]. This may be further investigated by applying stricter VIF thresholds, which was not addressed in the present study. Finally, information theory filters performed poorly, exhibiting both high FPR and low TPR, consistent with their omission of known biomarkers in the real outcome analysis.
Overall, a low FPR did not necessarily translate into better predictive performance, while higher TPR values were strongly associated with a higher AUC (Spearman’s $r=0.8$, $p<0.001$). This aligns with the intuitive expectation that strong generalization relies on identifying a sufficiently rich set of true predictors, rather than focusing solely on minimizing false positives.
Table 2 additionally compares computational efficiency across FS methods. Most filter and embedded methods required $\leq 10$ seconds per bootstrap iteration (and $\leq 1$ second for some). Wrapper methods and resampling-based approaches were substantially more demanding, with computation time approaching 1 minute per bootstrap iteration.
3.5 p.adjust filter achieved a more stable and predictive signature
The design of the PIONeeR study prioritised predictive performance, stability, and discovery capability (i.e., high TPR) in the evaluation of FS methods. A higher FPR was considered acceptable, as the cost of missing a true biomarker was greater than that of including redundant biomarkers. These considerations led to the identification of the p.adjust filter using the union of BH-adjusted p-values from logistic regression (adjusted for PD-L1) and the t-test at a 0.05 threshold as the top-performing method for predicting resistance to immunotherapy on PIONeeR data. The p.adjust demonstrated the highest optimism-corrected AUC (0.72) when fitted with the gradient boosting classifier. It also ranked among the most stable methods (S = 0.39). When benchmarked on semi-synthetic dataset, it achieved the highest TPR of 0.93 ± 0.06 at the expense of FPR of 0.1 ± 0.03.
Importantly, unlike standard LASSO, a baseline embedded FS method, and Shapicant, a top-performing representative of wrapper methods, all p.adjust-derived features were selected in $50\%$ of bootstrap samples, increasing the reliability of the resulting signature (Figure 5A). Additionally, we evaluated the stability of the p.adjust models at the individual patient level as measured by the instability index [44] – the proportion of bootstrap models that classify a given patient differently compared to the full model (Figure 5B). Ideally, the instability index should be elevated only for patients whose predicted probabilities lie close to the decision threshold of the full model. All models lacked this behavior. However, overall p.adjust improved patient-level stability: only 23.4% of patients had an instability index exceeding 0.1, compared to 29.9% for Shapicant and 32.0% for LASSO. At the more strict threshold of 0.3, this proportion decreased to 9.4% of patients for p.adjust, versus 12.4% for Shapicant and 9.9% for LASSO.
Although the signature selected by the p.adjust on the full dataset comprised 22 features, bootstrap-derived subsets included up to 72 features (Figure 5C). We investigated this by repeating the experiment using subsampling without replacement at 80% of the original sample size. Under subsampling, the inflation of the selected subset size disappeared, while stability increased and predictive performance was preserved. A similar reduction in selected subset size was observed for the majority of benchmarked methods under resampling without replacement. This suggests that the larger subset sizes observed with bootstrapping may result from over-representation of specific patient subclusters, increasing sample-specific signals that are weak in the full dataset. Therefore the sparsity instability of p.adjust appears to result from the bootstrapping procedure rather than the method itself.
Figure 5: Comparison of performance between Shapicant (representative wrapper method), LASSO (baseline embedded method), and p.adjust (the FS approach selected using ROOFS). A: Selection frequency of signature features. B: Instability index, defined as the proportion of bootstrap models leading to a different classification decision for a given patient compared to the full model [44]. C: Out-of-bag AUC ($\theta_{\text{OOB}}$; see Methods) as a function of selected subset size across bootstrap models.
In conclusion, conducting the FS benchmark using ROOFS and comparing the results obtained by multiple FS methods allowed us to improve signature performance and robustness compared to the baseline FS approach (LASSO) by identifying p.adjust as a method with optimal stability and AUC on the data with the real outcome and the best balance between TPR and FPR on the data with the simulated outcome.
4 Discussion
This paper introduced ROOFS, a Python framework designed to help researchers identify an FS method adapted to their predictive task. FS is a methodologically mature field, with extensive advances in algorithmic development [34], stability assessment [41], and false discovery control [24]. However, many of the theoretical developments have not been fully adopted in applied biomedical research, where practitioners often rely on a relatively narrow set of familiar methods. Moreover, FS is known to be data-dependent, with method performance varying across datasets and no single approach being universally optimal [28, 47, 16]. Nevertheless, FS methods are often applied with little consideration of their suitability for the specific dataset at hand. ROOFS addresses these gaps by integrating the evaluation of over 30 FS methods from different algorithmic families into a single pipeline that requires low to zero additional programming. To facilitate method comparison, ROOFS generates automatic reports summarizing FS performance metrics, as well as measures of individual feature robustness. A key element of ROOFS is its emphasis on stability – a property of FS that is frequently ignored in biomedical studies and in the development of new FS approaches despite its importance for reproducibility.
To our knowledge, no previously released packages combine multiple pre-implemented FS methods with the optimism correction framework for model evaluation. The closest existing tool is the R package pminternal (https://github.com/stephenrho/pminternal), which estimates optimism of the user-defined model functions, including those that contain an FS step. However, it does not provide any pre-implemented FS methods.
Currently, ROOFS is limited to classification tasks. Future work will focus on extending it to regression and survival tasks. Another limitation is that ROOFS requires considerable runtime to perform a full benchmark. While all implemented methods are scalable to moderately high-dimensional data, in datasets with thousands of features, it may be reasonable to reduce the runtime by avoiding wrapper and ensemble methods and limiting the number of classifiers used in downstream prediction.
We demonstrated the utility of ROOFS through application to the dataset from the PIONeeR clinical trial, focused on predicting resistance to anti-PD-(L)1 immunotherapy in advanced lung cancer. It enabled identification of the adjusted p-values filter – the union of a t-test and logistic regression controlled for PD-L1 at a 0.05 threshold – as an FS method offering the best stability, confidence in selected features, predictive performance and balance between TPR and FPR. The general findings of this study align with previous benchmarks where simple statistical filters outperformed more complicated ML approaches [23]. Therefore, we recommend starting the analysis of new datasets by benchmarking these methods, given their additional advantage of computational efficiency. ROOFS was also recently successfully applied to the analysis of the SChISM clinical study (Size CfDNA Immunotherapy Signature Monitoring, NCT05083494). It allowed to identify a nine-variables signature derived from fragmentomics circulating free DNA data, with enhanced predictive power of resistance (AUC = 0.89 ± 0.033, PPV = 0.69 ± 0.075) [40].
To conclude, the choice of an FS method and, therefore, the resulting subset of features or "signature", should be driven by the dataset’s characteristics and study goals. It should consider study-specific acceptable trade-off between model stability, performance, and relative costs of false positives and false negatives. ROOFS can serve as a practical tool for researchers working in settings where reproducibility is crucial. By enabling comprehensive benchmarking, ROOFS and similar tools have the potential to improve the reproducibility of biomarker discoveries and increase the translational value of clinical predictive models.
Key points
  • There is a significant disconnect between the rich methodological development of feature selection (FS) and its use in applied biomedical research. Practitioners frequently rely on a limited subset of FS approaches and do not use comprehensive evaluation procedures to assess the suitability of the method for a given dataset.
  • FS stability, defined as the consistency of selected features across data resamples, is essential for reproducibility and clinical translation of biomarker signatures, yet it often remains unreported in biomedical studies.
  • We developed a Python package, ROOFS, designed to benchmark multiple FS methods on user-provided data with emphasis on method stability, individual feature robustness, predictive performance, and error rates in a separate analysis with a simulated outcome.
  • We demonstrated ROOFS utility by applying it to a real-world problem: predicting resistance to immunotherapy in lung cancer patients from the PIONeeR clinical trial. This comparative study identified a statistical filter based on the union of Benjamini-Hochberg adjusted p-values from t-test and logistic regression (adjusted for PD-L1) at a 0.05 threshold as the optimal FS method across multiple evaluation metrics.
Data availability statement
The PIONeeR data underlying this article cannot be shared publicly due to patient privacy reasons. Two publicly available datasets are provided in the package for tutorial and testing [9, 39, 40], as well as one synthetic dataset.
Funding statement
This work was supported by the French National Research Agency (ANR) under the France 2030 investment plan through grants ANR-17-RHUS-0007 and ANR-22-PESN-0017, the latter awarded within the DIGPHAT project.
This work was supported by a partnership of Aix-Marseille Université (AMU), Assistance Publique Hôpitaux de Marseille (APHM), Centre National de La Recherche Scientifique (CNRS), Institut National de la Santé et de la Recherche Médicale (INSERM), Centre Léon Bérard (CLB), Institut Paoli Calmettes (IPC), Gustave Roussy (GR), AstraZeneca (AZ), Veracyte (VERA), Innate Pharma (IPH) & ImCheck Therapeutics (ICT), and initiated by Marseille Immunopole.
Acknowledgements
The authors gratefully acknowledge the support of the APHM, which sponsored the PIONeeR clinical study. Its role was to control the appropriateness of ethical and legal considerations for all centers and to perform the monitoring of the consents signed and the clinical data recorded and coded as part of the study. The authors are grateful to all the patients and their families, as well as all the investigators, for their participation in the study.
The authors are grateful to Linh Nguyen Phuong, PhD, and Mohamed Boussena for helpful comments.
Supplementary table S1: Benchmarked FS methods
Method Description Reference
Embedded methods
LASSO Logistic regression with the L1 penalty. [49]
Adaptive LASSO LASSO variant that applies adaptive weights to each coefficient, allowing to penalize unimportant variables more than important ones. [55]
[…]
Ensemble embedded methods
Bolasso Modified bootstrap LASSO. Selects features with a bootstrap frequency of selection of $\geq 0.5$ instead of 1. Our modification of [1]
Stability selection LASSO-based subsampling method that performs repeated fits across a range of regularization parameters and selects features selected with a frequency above a specified threshold. [37]
Stabl LASSO Recent extension of stability selection that injects artificial noise variables to enable data-driven choice of the selection frequency threshold. [24]
RENT Repeated Elastic Net Technique, a resampling-based method that selects features that satisfy three criteria: a sufficiently high selection frequency across resampled models, consistent coefficient sign across models, and a sufficiently large ratio of the mean coefficient to its standard error. [29]
Block HSIC LASSO Method that the Hilbert-Schmidt Independence Criterion (HSIC) to identify non-linear dependencies between features and the outcome while penalizing redundant features. To reduce memory complexity for the case of high-dimensional data, the method partitions samples into smaller blocks. The final bagging block HSIC estimator is computed by averaging HSIC values across multiple random permutations of the data partitions. [13]
Filter methods
CIFE Iteratively selects features with the highest Conditional Infomax Feature Extraction (CIFE) score until the user-specified set size is reached. For each feature $X_{k}$: $\text{CIFE}(X_{k})=I(X_{k};Y)-\sum_{X_{j}\in\mathcal{S}}I(X_{j};X_{k})+\sum_{X_{j}\in\mathcal{S}}I(X_{j};X_{k}\mid Y)$ The first term maximizes mutual information (MI) between the feature and the outcome, the second penalizes redundancy with previously selected features, and the third rewards conditional redundancy given the outcome. [34, 36]
CMIM Iteratively selects features with the highest Conditional Mutual Information Maximization (CMIM) score until the user-specified set size is reached. For each feature $X_{k}$: $\text{CMIM}(X_{k})=\text{min}_{X_{j}\in\mathcal{S}}[I(X_{k};Y\mid X_{j})]$ The CMIM score maximizes the minimal conditional MI between the candidate feature $X_{k}$ and the outcome $Y$, with the minimum value chosen across all previously selected features $X_{j}\in\mathcal{S}$. [34, 19]
DISR Iteratively selects features with the highest Double Input Symmetrical Relevance (DISR) score until the user-specified set size is reached. For each feature $X_{k}$: $\text{DISR}(X_{k})=\sum_{X_{j}\in F}\frac{I(X_{j}X_{k};Y)}{H(X_{j}X_{k}Y)}$, where $I(X_{j}X_{k};Y)=I(X_{k};Y)+I(X_{j};Y\mid X_{k})$, and $H(X_{j}X_{k}Y)=H(X_{k})+H(X_{k}|X_{j})+H(Y|X_{k})-I(Y;X_{j}|X_{k})$. [34, 38]
[…]
ReliefF Ranks features by a distance-based score and selects top-k features. For each sample, the algorithm penalizes the features with different values in the nearest neighbors (5 by default) of the same class and rewards the features with different values for the nearest neighbors (similarly, 5 by default) of the opposite class. [34, 46]
Fisher score Ranks features by the ratio of between-class variance to within-class variance and selects top-k features. [34]
Gini index Ranks features by their minimum Gini impurity (computed by testing all possible binary splits) and selects top-k features. [34]
t-score Ranks features by their two-sample t-statistic and selects top-k features. [34]
mRMR Iteratively adds features with the highest relevance-redundancy ratio until the desired number of features is reached. Relevance is measured by the ANOVA F-value, redundancy is iteratively updated as the mean correlation with previously selected features.
Hierarchical clustering Performs hierarchical clustering of features based on Spearman correlation coefficients, with the number of clusters equal to the desired number of features to be selected. From each cluster, the feature with the highest point biserial correlation with the outcome is selected. Our modification of [43]
p.adjust t-test + logistic .05 Union-based filter that combines features selected by t-test and logistic regression (controlling for user-specified covariates; PD-L1 in the case of PIONeeR data), where p-values from each test were adjusted using the Benjamini-Hochberg procedure with a 0.05 threshold. Our modification of p-value filtering
Wrapper methods
LR-RFE Recursive feature elimination that iteratively removes 10% of features with the lowest absolute coefficients from a logistic regression model with the L2 penalty until the pre-specified number of features is reached. [21]
RF-RFE Recursive feature elimination that iteratively removes 10% of the least important feature of a random forest model and selects a subset with the best AUC in cross-validation. [30]
Forward RF Recursive forward selection that iteratively adds the most important feature of a random forest model until a pre-specified set size is reached and selects a subset with the best AUC in cross-validation. [50]
Shapicant Wrapper that uses SHAP importances together with outcome permutation. The criterion of selection is the threshold on p-values from testing whether each feature’s SHAP value under the real outcome differs from that under the permuted outcome. In ROOFS, the default procedure begins with a threshold of 0.1 and iteratively decreases it, stopping after three iterations or once the number of selected features $k$ falls below the user-specified max_num_selected. [8]
Controls
Random FS Selects a feature subset randomly. For the benchmark on PIONeeR data, subset sizes $k$ were randomly sampled from the range 1-20 for the real dataset. For the semi-synthetic dataset, subset size was either set to $k=19$ (a fixed size corresponding to the number of predictors in the true model) or sampled randomly from [1, $p$].
Full model Model trained with all features.
Supplementary table S2: ML classifiers used in FS benchmark on PIONeeR dataset
Category Classifier Hyperparametersa
Tree-based BaggingClassifier base_estimator=DecisionTree, n_estimators=200
max_depth=2, min_samples_leaf=40, max_features=0.7
ExtraTreesClassifier n_estimators=200, max_depth=2
min_samples_leaf=40, max_features=0.7
RandomForestClassifier n_estimators=200, max_depth=2
min_samples_leaf=40, max_features=0.7
Boosting GradientBoostingClassifier max_depth=2, min_samples_leaf=40
max_features=0.7, subsample=0.7
HistGradientBoostingClassifier max_depth=2, min_samples_leaf=40, max_features=0.7
XGBClassifier max_depth=2, n_jobs=-1, other parameters at default
Linear LogisticRegression penalty=None, n_jobs=-1
LinearDiscriminantAnalysis Default hyperparameters
Other KNeighborsClassifier Default
MLPClassifier Default
SVC probability=True, other parameters at default
Ensemble StackingClassifierb cv=5, stack_method=predict_proba
Base estimators: LogisticRegression, RandomForest,
SVC, GradientBoosting
Final estimator: LogisticRegression
a All classifiers use scikit-learn default parameters unless specified.
b StackingClassifier’s base estimators use the same hyperparameters as the individual classifiers.
Supplementary figure S1: Performance of all benchmarked FS methods (top: FS methods with fixed, user-defined subset size; bottom: FS methods with subset size based on algorithm thresholds). Blue: embedded methods; violet-blue: resampling-based ensemble methods; green: filters; rose: wrappers.A, C: False omission and false discovery rates in the experiment on the semi-synthetic PIONeeR dataset with a simulated outcome. The false omission rate represents the proportion of true features that were incorrectly not selected (false negatives) among all non-selected features, while the false discovery rate represents the proportion of incorrectly selected features (false positives) relative to the size of the signature derived by an FS algorithm. Error bars represent one-half of the standard deviation (SD). B, D: Stability and predictive performance (AUC) in the experiment on the real PIONeeR dataset.